# Posterior Ramifications of Prior Dependence Structures

### 11 Dec 2023 v1

@@ 4.1 Convergence of the Posterior Mode
We can factor out the $f_{j}(\theta_{j})$ term because $p_{1}(\boldsymbol{\theta})$ and $p_{2}(\boldsymbol{\theta})$ are defined using the same marginals. The priors were also defined such that $p_{1}(\boldsymbol{\theta})>0$ and $p_{2}(\boldsymbol{\theta})>0$ for all $\boldsymbol{\theta}\in\boldsymbol{\Theta}$. Therefore, $f_{j}(\theta_{j})$ must be positive, and the partial derivative in (7) with respect to $\theta_{j}$ is 0 if and only if the partial derivative with respect to $u_{j}$ is 0. This correspondence ensures the results from Theorem 2 generalize over different specifications for the marginal distributions. We prove both parts of Theorem 2 below.
Proof of Theorem 2(a). The posterior mode $\boldsymbol{\theta}^{{}_{(k)}}$ minimizes $-\text{log}(p_{k}(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}}))$ for $k=1,2$. We have that
$\begin{split}&-\text{log}(p_{2}(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}}))\\ &=-l(\boldsymbol{\theta};\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}}))-\sum_{j=1}^{d}\text{log}(f_{j}(\theta_{j}))-\text{log}(c_{1}(\boldsymbol{u}))+\\ &~{}~{}~{}~{}~{}~{}(\text{log}(c_{1}(\boldsymbol{u}))-\text{log}(c_{2}(\boldsymbol{u}))),\end{split}$ (8)
where $l(\boldsymbol{\theta};\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}}))$ is the log-likelihood function. The first three terms in (8) are equal to $-\text{log}(p_{1}(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}}))$. When $\boldsymbol{\theta}=\boldsymbol{\theta}^{{}_{(1)}}$, it follows that $-\nabla_{\boldsymbol{\theta}}\text{log}(p_{1}(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}}))=0$ and $-\nabla_{\boldsymbol{\theta}}\text{log}(p_{2}(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}}))=-\nabla_{\boldsymbol{\theta}}\big{[}\text{log}(c_{1}(\boldsymbol{u}))-\text{log}(c_{2}(\boldsymbol{u}))\big{]}$. For large sample sizes $n$, the conditions for Theorem 2 ensure that the log-posterior $-\text{log}(p_{2}(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}}))$ converges to its quadratic approximation about $\boldsymbol{\theta}=\boldsymbol{\theta}_{0}$. We use $\boldsymbol{\theta}^{{}_{(1)}}$ as a starting point for the Newton–Raphson method (Nocedal and Wright, 2006) with $-\text{log}(p_{2}(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}}))$ to find $\boldsymbol{\theta}^{{}_{(2)}}$. Because the quadratic approximation is appropriate for large samples, only one iteration of the Newton–Raphson method is required to approximate $\boldsymbol{\theta}^{{}_{(2)}}$. We let $\mathcal{J}(\cdot)$ be the observed information. It follows that
$\boldsymbol{\theta}^{{}_{(2)}}\approx\boldsymbol{\theta}^{{}_{(1)}}-\mathcal{J}(\boldsymbol{\theta}^{{}_{(1)}})^{-1}\nabla_{\boldsymbol{\theta}=\boldsymbol{\theta}^{{}_{(1)}}}\big{[}\text{log}(c_{1}(\boldsymbol{u}))-\text{log}(c_{2}(\boldsymbol{u}))\big{]}.$ (9)
The conditions for Theorem 2 also guarantee that the posterior mode $\boldsymbol{\theta}^{{}_{(1)}}$ is approximately equal to the approximately normal MLE $\hat{\boldsymbol{\theta}}^{{}_{(1)}}$ for large samples $\boldsymbol{y}^{{}_{(n)}}$. Because the MLE is consistent, $\mathcal{J}(\boldsymbol{\theta}^{{}_{(1)}})^{-1}\approx\mathcal{J}(\boldsymbol{\theta}_{0})^{-1}$ for sufficiently large samples. By (7), we have that $\nabla_{\boldsymbol{\theta}=\boldsymbol{\theta}_{0}}\big{[}\text{log}(c_{1}(\boldsymbol{u}))-\text{log}(c_{2}(\boldsymbol{u}))\big{]}\neq\boldsymbol{0}$, so $\text{log}(c_{1}(\boldsymbol{u}))-\text{log}(c_{2}(\boldsymbol{u}))$ can be approximated by a plane (with common gradient) in a neighbourhood of $\boldsymbol{\theta}_{0}$. It follows that $\mathcal{J}(\boldsymbol{\theta}^{{}_{(1)}})^{-1}\nabla_{\boldsymbol{\theta}=\boldsymbol{\theta}^{{}_{(1)}}}\big{[}\text{log}(c_{1}(\boldsymbol{u}))-\text{log}(c_{2}(\boldsymbol{u}))\big{]}$ will be roughly constant for all large samples $\boldsymbol{y}^{{}_{(n)}}$ generated from $m(\hskip 1.0pt\cdot\hskip 1.0pt|\hskip 1.0pt\boldsymbol{\theta}_{0})$. For an arbitrary large sample $\boldsymbol{y}^{{}_{(n)}}$, we have that $\boldsymbol{\theta}^{{}_{(2)}}\approx\boldsymbol{\theta}^{{}_{(1)}}+\boldsymbol{b}$ for some constant $\boldsymbol{b}$ that does not depend on $\boldsymbol{y}^{{}_{(n)}}$ by (9). Since the posterior concentrates around $\boldsymbol{\theta}_{0}$ as the sample size $n$ increases, this common perturbation $\boldsymbol{b}$ will decrease in magnitude. $\boldsymbol{\theta}^{{}_{(1)}}$ is approximately normally distributed about $\boldsymbol{\theta}_{0}$ for large samples, so these small perturbations will shift $\boldsymbol{\theta}^{{}_{(2)}}\approx\boldsymbol{\theta}^{{}_{(1)}}+\boldsymbol{b}$ closer to $\boldsymbol{\theta}_{0}$ with probability of roughly 0.5 due to the symmetry of the normal distribution.
Proof of Theorem 2(b). The results from (8) and (9) hold true as in part (a). Because $\boldsymbol{u}_{0}$ is a local minimum of $\text{log}(c_{1}(\boldsymbol{u}))-\text{log}(c_{2}(\boldsymbol{u}))$, the local linear approximation of this function is not serviceable at $\boldsymbol{\theta}=\boldsymbol{\theta}_{0}$. However, this implies that $-\nabla_{\boldsymbol{\theta}=\boldsymbol{\theta}^{{}_{(1)}}}\big{[}\text{log}(c_{1}(\boldsymbol{u}))-\text{log}(c_{2}(\boldsymbol{u}))\big{]}$ should be directed toward $\boldsymbol{\theta}_{0}$ for large samples $\boldsymbol{y}^{{}_{(n)}}$ when the quadratic approximation to the log-posterior is appropriate and $\boldsymbol{\theta}^{{}_{(1)}}\approx\boldsymbol{\theta}_{0}$. We let the orthonormal basis $\mathcal{B}$ be composed of the eigenvectors of $\mathcal{J}(\boldsymbol{\theta}^{{}_{(1)}})^{-1}$. Since $\mathcal{J}(\boldsymbol{\theta}^{{}_{(1)}})^{-1}$ is a positive definite matrix, the angle between $-\nabla_{\boldsymbol{\theta}=\boldsymbol{\theta}^{{}_{(1)}}}\big{[}\text{log}(c_{1}(\boldsymbol{u}))-\text{log}(c_{2}(\boldsymbol{u}))\big{]}$ and $-\mathcal{J}(\boldsymbol{\theta}^{{}_{(1)}})^{-1}\nabla_{\boldsymbol{\theta}=\boldsymbol{\theta}^{{}_{(1)}}}\big{[}\text{log}(c_{1}(\boldsymbol{u}))-\text{log}(c_{2}(\boldsymbol{u}))\big{]}$ will be acute. With respect to $\mathcal{B}$, the perturbation from $\boldsymbol{\theta}^{{}_{(1)}}$ to $\boldsymbol{\theta}^{{}_{(2)}}$ induced by the Newton-Raphson method is then directed in the same orthant of $\boldsymbol{\Theta}\subseteq\mathbb{R}^{d}$ that contains $\boldsymbol{\theta}_{0}$. It follows that for large samples, $\boldsymbol{\theta}^{{}_{(2)}}$ cannot be further from $\boldsymbol{\theta}_{0}$ than $\boldsymbol{\theta}^{{}_{(1)}}$ due to a perturbation in the wrong direction. However, $\boldsymbol{\theta}^{{}_{(2)}}$ could still be further from $\boldsymbol{\theta}_{0}$ than $\boldsymbol{\theta}^{{}_{(1)}}$ if the magnitude of the perturbation is too large. We argue that this cannot occur for an arbitrary large sample $\boldsymbol{y}^{{}_{(n)}}$ because both $p_{1}(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}}))$ and $p_{2}(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}}))$ will concentrate around $\boldsymbol{\theta}_{0}$ and $\boldsymbol{\theta}_{0}$ maximizes $\text{log}(c_{2}(\boldsymbol{u}))-\text{log}(c_{1}(\boldsymbol{u}))$ in a neighbourhood of this fixed point. If this perturbation is too large for a given sample $\boldsymbol{y}^{{}_{(n)}}$, this behaviour cannot persist as $n\rightarrow\infty$. For large samples, these small perturbations by the Newton-Raphson method will therefore shift $\boldsymbol{\theta}^{{}_{(2)}}$ closer to $\boldsymbol{\theta}_{0}$ with probability approaching 1.
@@
We note that if $\boldsymbol{u}_{0}$ is a local minimum of $\text{log}(c_{2}(\boldsymbol{u}))-\text{log}(c_{1}(\boldsymbol{u}))$, then $\lim_{n\rightarrow\infty}Pr(\mathcal{D}_{2}\leq\mathcal{D}_{1})=0$. This follows directly from part (b) of Theorem 2 by switching the labels on $c_{1}(\cdot)$ and $c_{2}(\cdot)$. The case where $\boldsymbol{u}_{0}$ is a saddle point of $\text{log}(c_{2}(\boldsymbol{u}))-\text{log}(c_{1}(\boldsymbol{u}))$ is excluded from both parts (a) and (b). In that case, $Pr(\mathcal{D}_{2}\leq\mathcal{D}_{1})$ may converge to a constant that is not 0.5 or 1. We explore the results from Theorem 2 via simulation and explain their practical implications in Section 4.2.
@@ 4.2 Practical Implications
Here, we conduct simulations to consider Theorem 2 in practice. To do so, we consider an example adapted from Michimae and Emura (2022), who recently suggested that joint priors based on Archimedean copulas lead to more accurate and concentrated posterior distributions in the context of Bayesian ridge regression. These recommendations are inconsistent with the results of Theorem 2. The accuracy and concentration of the posterior of $\boldsymbol{\theta}$ were considered via the total mean absolute error between the posterior median and a fixed parameter value $\boldsymbol{\theta}_{0}$. Michimae and Emura’s (2022) numerical studies showed that the posterior was more concentrated around $\boldsymbol{\theta}_{0}$ when the marginal priors for the regression coefficients were joined using (Archimedean) Clayton and Gumbel copulas (Nelsen, 2006) instead of more standard Gaussian copulas. They considered a ridge regression model with three regression coefficients in the presence of multicollinearity, where the parameters of the relevant prior copulas were random variables specified using a hierarchical framework. The simplified example for our numerical study adapts aspects of their model for illustrative purposes.
@@
where $\varepsilon_{i}\sim\mathcal{N}(0,5)$ independently for $i=1,...,n$. The assumptions that the linear equation has an intercept of zero and the error terms have known variance reduce the dimensionality of the problem for illustration. That is, $\boldsymbol{\theta}=(\beta_{1},\beta_{2})$. We specify standard normal marginal priors for both $\beta_{1}$ and $\beta_{2}$.
We join these marginal priors with two prior copulas in this numerical study: $c_{1}(\boldsymbol{u})=1$ for $\boldsymbol{u}\in(0,1)^{2}$ corresponds to the independence copula and $c_{2}(\boldsymbol{u})$ corresponds to a two-dimensional $t$-copula with $\nu=4$ degrees of freedom and a diagonal correlation matrix $\boldsymbol{R}$. We select the first copula because it is often assumed that the regression coefficients are independent a priori in Bayesian regression models. The corresponding joint prior for $\beta_{1}$ and $\beta_{2}$ is therefore a standard bivariate normal distribution with diagonal $\boldsymbol{R}$. The choice for the second copula is motivated by $c_{2}(\boldsymbol{u})$ having a local maximum and saddle points to illustrate the results from Theorem 2. Figure 3 visualizes the logarithm of this copula density function. The selected $t$-copula does not accommodate strong negative or positive dependence between $\beta_{1}$ and $\beta_{2}$, but it reflects a greater likelihood of observing extreme values for both $\beta_{1}$ and $\beta_{2}$ relative to their marginal priors. For instance, this might occur if both marginal priors were misspecified.
@@
We consider six values for $\boldsymbol{\theta}_{0}$ to define the data generation process for $\boldsymbol{Y}^{{}_{(n)}}$. These six values are meant to illustrate the convergence of the posterior mode in a variety of settings. Because $\boldsymbol{u}_{0}=(F_{1}(\theta_{1,0},F_{2}(\theta_{2,0})))$, we can readily convert between $\boldsymbol{\theta}_{0}$ and $\boldsymbol{u}_{0}$ given the specified $\mathcal{N}(0,1)$ marginals. The marginal priors for $\beta_{1}$ and $\beta_{2}$ were chosen to be rather informative so that we observe a range of behaviour for the settings corresponding to part $(a)$ of Theorem 2.
For each $\boldsymbol{\theta}_{0}$ value, we generated 10000 samples of size $n$ for various sample sizes between 5 and $10^{5}$. Each observation was simulated independently as $y_{i}=\theta_{1,0}x_{i1}+\theta_{2,0}x_{i2}+\varepsilon_{i}$, where $(x_{i1},x_{i2})\sim\mathcal{N}(\boldsymbol{0},I_{2})$ and $\varepsilon_{i}\sim\mathcal{N}(0,5)$ for $i=1,...,n$. For each of these 10000 samples, we found both posterior modes $\tilde{\boldsymbol{\theta}}^{{}_{(1)}}$ and $\tilde{\boldsymbol{\theta}}^{{}_{(2)}}$. We then estimated $Pr(\mathcal{D}_{2}\leq\mathcal{D}_{1})$ as the proportion of samples for which $\tilde{\boldsymbol{\theta}}^{{}_{(2)}}$, corresponding to the $t$-copula, was closer to $\boldsymbol{\theta}_{0}$ than $\tilde{\boldsymbol{\theta}}^{{}_{(1)}}$. To consider the practical impact of Theorem 2, we also computed the mean absolute difference between $\mathcal{D}_{2}$ and $\mathcal{D}_{1}$ for each sample size $n$. Figure 4 visualizes these results for each of the six scenarios, which we now describe.
@@
We now contrast our conclusions from this numerical study with Michimae and Emura’s (2022) recommendations. First, the Clayton and Gumbel copulas used by Michimae and Emura (2022) only accommodate positive prior dependence between $\beta_{1}$ and $\beta_{2}$. If the covariates $x_{1}$ and $x_{2}$ are positively correlated, then the off-diagonal components in the inverse Fisher information matrix $\mathcal{I}(\tilde{\boldsymbol{\theta}})^{-1}$ evaluated at the posterior mode $\tilde{\boldsymbol{\theta}}$ will be negative for sufficiently large $n$. In this case, $\beta_{1}$ and $\beta_{2}$ will be unable to retain a postive dependence structure a posteriori. However, this does not constitute chronic rejection as defined in Section 2 because it is possible for $\beta_{1}$ and $\beta_{2}$ to retain their positive dependence a posteriori if $x_{1}$ and $x_{2}$ are negatively correlated. This example illustrates that the conditions for Theorem 1 are sufficient in that the prior dependence structure cannot be retained as data are observed when these conditions are satisfied. However, it is not guaranteed that the prior dependence structure will be retained when the conditions for Theorem 1 are not satisfied.

### 23 Mar 2024 v2

@@ 4.1 Convergence of the Posterior Mode
We can factor out the $f_{j}(\theta_{j})$ term because $p_{1}(\boldsymbol{\theta})$ and $p_{2}(\boldsymbol{\theta})$ are defined using the same marginals. The priors were also defined such that $p_{1}(\boldsymbol{\theta})>0$ and $p_{2}(\boldsymbol{\theta})>0$ for all $\boldsymbol{\theta}\in\boldsymbol{\Theta}$. Therefore, $f_{j}(\theta_{j})$ must be positive, and the partial derivative in (8) with respect to $\theta_{j}$ is 0 if and only if the partial derivative with respect to $u_{j}$ is 0. This correspondence ensures the results from Theorem 2 generalize over different specifications for the marginal distributions.
@@
We prove Theorem 2 in Appendix C of the online supplement. We note that if $\boldsymbol{u}_{0}$ is a local minimum of $\text{log}(c_{2}(\boldsymbol{u}))-\text{log}(c_{1}(\boldsymbol{u}))$, then $\lim_{n\rightarrow\infty}Pr(\mathcal{D}_{2}\leq\mathcal{D}_{1})=0$. This follows directly from part (b) of Theorem 2 by switching the labels on $c_{1}(\cdot)$ and $c_{2}(\cdot)$. The case where $\boldsymbol{u}_{0}$ is a saddle point of $\text{log}(c_{2}(\boldsymbol{u}))-\text{log}(c_{1}(\boldsymbol{u}))$ is excluded from both parts (a) and (b). In that case, $Pr(\mathcal{D}_{2}\leq\mathcal{D}_{1})$ may converge to a constant that is not 0.5 or 1. We explore the results from Theorem 2 via simulation and explain their practical implications in Section 4.2.
@@ 4.2 Practical Implications
Here, we conduct simulations to consider Theorem 2 in practice. To do so, we consider an example adapted from Michimae and Emura (2022) since their recommendations are inconsistent with the results of Theorem 2. They considered a ridge regression model with three regression coefficients in the presence of multicollinearity, where the parameters of the relevant prior copulas were random variables specified using a hierarchical framework. The simplified example for our numerical study adapts aspects of their model for illustrative purposes. Note that we contrast our results with Michimae and Emura’s (2022) findings in Section 4.3.
@@
where $\varepsilon_{i}\sim\mathcal{N}(0,5)$ independently for $i=1,\dots,n$. The assumptions that the linear equation has an intercept of zero and the error terms have known variance reduce the dimensionality of the problem for illustration. That is, $\boldsymbol{\theta}=(\beta_{1},\beta_{2})$. We specify standard normal marginal priors for both $\beta_{1}$ and $\beta_{2}$.
We join these marginal priors with two prior copulas in this numerical study: $c_{1}(\boldsymbol{u})=1$ for $\boldsymbol{u}\in[0,1]^{2}$ corresponds to the independence copula and $c_{2}(\boldsymbol{u})$ corresponds to a two-dimensional $t$-copula with $\nu=4$ degrees of freedom and a diagonal correlation matrix $\boldsymbol{R}$. We select the first copula because it is often assumed that the regression coefficients are independent a priori in Bayesian regression models. The corresponding joint prior for $\beta_{1}$ and $\beta_{2}$ is therefore a standard bivariate normal distribution with diagonal $\boldsymbol{R}$. The choice for the second copula is motivated by $c_{2}(\boldsymbol{u})$ having a local maximum and saddle points to illustrate the results from Theorem 2. Figure 4 visualizes the logarithm of this copula density function. The selected $t$-copula does not accommodate strong negative or positive dependence between $\beta_{1}$ and $\beta_{2}$, but it reflects a greater likelihood of observing extreme values for both $\beta_{1}$ and $\beta_{2}$ relative to their marginal priors. For instance, this might occur if both marginal priors were misspecified.
@@
We consider six values for $\boldsymbol{\theta}_{0}$ to define the data generation process for $\boldsymbol{Y}^{{}_{(n)}}$. These six values are meant to illustrate the convergence of the posterior mode in a variety of settings. Because $\boldsymbol{u}_{0}=(F_{1}(\theta_{1,0}),F_{2}(\theta_{2,0}))$, we can readily convert between $\boldsymbol{\theta}_{0}$ and $\boldsymbol{u}_{0}$ given the specified $\mathcal{N}(0,1)$ marginals. The marginal priors for $\beta_{1}$ and $\beta_{2}$ were chosen to be rather informative so that we observe a range of behaviour for the settings corresponding to part $(a)$ of Theorem 2.
For each $\boldsymbol{\theta}_{0}$ value, we generated 10000 samples of size $n$ for various sample sizes between 5 and $10^{5}$. Each observation was simulated independently as $y_{i}=\beta_{1,0}x_{i1}+\beta_{2,0}x_{i2}+\varepsilon_{i}$, where $(x_{i1},x_{i2})\sim\mathcal{N}(\boldsymbol{0},I_{2})$ and $\varepsilon_{i}\sim\mathcal{N}(0,5)$ for $i=1,\dots,n$. For each of these 10000 samples, we found both posterior modes $\tilde{\boldsymbol{\theta}}^{{}_{(1)}}$ and $\tilde{\boldsymbol{\theta}}^{{}_{(2)}}$ using calculus. We then estimated $Pr(\mathcal{D}_{2}\leq\mathcal{D}_{1})$ as the proportion of samples for which $\tilde{\boldsymbol{\theta}}^{{}_{(2)}}$, corresponding to the $t$-copula, was closer to $\boldsymbol{\theta}_{0}$ than $\tilde{\boldsymbol{\theta}}^{{}_{(1)}}$. To consider the practical impact of Theorem 2, we also computed the mean absolute difference between $\mathcal{D}_{2}$ and $\mathcal{D}_{1}$ for each sample size $n$. Figure 5 visualizes these results for each of the six scenarios, which we now describe.
@@
Table 2: Empirical coverage of 95% HPD sets for $\boldsymbol{\theta}=(\beta_{1},\beta_{2})$ across 10000 posteriors defined using both prior copulas.
Case for Independence Copula
$n$ 1 2 3 4 5 6
$10^{1}$ 0.9928 0.9916 0.9685 0.8723 0.9465 0.0023
$10^{2}$ 0.9574 0.9575 0.9558 0.9387 0.9479 0.5698
$10^{3}$ 0.9510 0.9550 0.9515 0.9509 0.9504 0.9181
$10^{4}$ 0.9519 0.9525 0.9517 0.9546 0.9542 0.9452
$10^{5}$ 0.9546 0.9501 0.9507 0.9523 0.9473 0.9508
Case for $t$-Copula
$n$ 1 2 3 4 5 6
$10^{1}$ 0.9939 0.9938 0.9683 0.7769 0.9465 0.0002
$10^{2}$ 0.9583 0.9603 0.9570 0.9234 0.9503 0.5148
$10^{3}$ 0.9509 0.9558 0.9532 0.9481 0.9507 0.9162
$10^{4}$ 0.9515 0.9532 0.9524 0.9512 0.9553 0.9451
$10^{5}$ 0.9539 0.9503 0.9518 0.9516 0.9479 0.9524
@@
For each case and prior combination, we also estimated empirical coverage as the proportion of the 10000 posteriors for which the 95% HPD set included the parameter value $\boldsymbol{\theta}_{0}$ when $n=\{10^{1},10^{2},10^{3},10^{4},10^{5}\}$. The HPD sets were again estimated using two-dimensional kernel density estimation on posterior draws obtained via sampling-resampling methods. The empirical coverage results are summarized for both prior copulas in Table 2.
Per Table 2, the empirical coverage is similar for both prior copulas at all samples sizes considered in cases 1, 2, 3, and 5. For small sample sizes, the empirical coverage exceeds the nominal value of 95% in cases 1, 2, and 3. Since this trend is observed for both prior copulas, it is caused by informative nature of the relatively well-specified marginal priors for $\beta_{1}$ and $\beta_{2}$. These marginal priors are misspecified in cases 4 and 6, so the empirical coverage is less than 95% for both prior copulas when the sample size is small. The empirical coverage is better when using the independence copula for small $n$ in both cases – even though $\tilde{\boldsymbol{\theta}}^{{}_{(2)}}$ was generally closer to $\boldsymbol{\theta}_{0}$ than $\tilde{\boldsymbol{\theta}}^{{}_{(1)}}$ for case 6 when $n$ was small in Figure 5.
4.3 Connections to Previous Contributions